darkknight:

f(x) = x/(x-9) find (f^-1) ' (7/16) (the derivative of the inverse function)

darkknight:

ok why is this wrong, first i find the inverse y= x/(x-9) x = y/(y-9) (y-9)x = y yx-9x=y yx-y = 9x y(x-1)=9x y = 9x/(x-1) f^-1(x) = 9x/(x-1)

darkknight:

found the inverse, now i plug into the formula \[1/(f'(f^-1(x))\] So we find the derivative of this inverse, now we use quotient rule (x-1)(9)-(9x)(1))/(x-1)^2 which simplifies down to -9/(x-1)^2 now the derivative of the inverse is 1/ all that so (x-1)^2/-9 is the derivative of the inverse I plugged in 7/16 into the equation and got -9/256. What did I do wrong???

darkknight:

@dude

darkknight:

@imqwerty @jimthompson5910 @jhonyy9

imnotsmartlol:

what is this math im scared to move on

jimthompson5910:

hint: If \(f(x) = \frac{x}{x-9}\), then \( f'(x) = -\frac{9}{(x-9)^2}\)

darkknight:

Yeah thats what I did, according to my formula the derivative of an inverse function is \[d/dx(f^-1(x)) = 1/f'(f^-1(x))\]

darkknight:

So since its 1 over that part I just took (x-9)^2/-9 and plugged in 7/16 for x

jimthompson5910:

You need to compute \(f^{-1}\left(\frac{7}{16}\right)\) first \(f^{-1}(x) = \frac{9x}{x-1}\) \(f^{-1}\left(\frac{7}{16}\right) = \frac{9\left(\frac{7}{16}\right)}{\frac{7}{16}-1}\) \(f^{-1}\left(\frac{7}{16}\right) = m\) The goal is to find \(\frac{1}{f'(m)}\)

darkknight:

does it not work the way I do it? where I find the inverse and then solve for the derivative of the inverse and lastly plug in the number? that's what my teacher taught so that's y i am confused. He has an example for a function f(x) = x^2 the inverse is \[f^-1(x) = \sqrt{x}\] and then find derivative you get 1/2x^(-1/2) and 1 / by that so you get 1/(2sqrt(x))

darkknight:

OMG me big bot. I get what you are saying now my b. I understood. Ty jimthompson

jimthompson5910:

If you follow that method, then you won't need the formula \(\frac{d}{dx}[f^{-1}(x)] = \frac{1}{f'(f^{-1}(x))}\)

darkknight:

yep, i get it now thx jimthompson, i have like 2 more questions if u don't mind.

jimthompson5910:

go ahead

darkknight:

yeah so what happened is my teacher taught 2 different methods, i did some weird thing and combined both of them idk why